# Predicting the Distribution of Treatment Effects: A Covariate-Adjustment Approach

### v2

@@ Abstract
Important questions for impact evaluation require knowledge not only of average effects, but of the distribution of treatment effects. What proportion of people are harmed? Does a policy help many by a little? Or a few by a lot? The inability to observe individual counterfactuals makes these empirical questions challenging. I propose an approach to inference on points of the distribution of treatment effects by incorporating predicted counterfactuals through covariate adjustment. I show that finite-sample inference is valid under weak assumptions, for example, when data come from a Randomized Controlled Trial (RCT), and that large-sample inference is asymptotically exact under suitable conditions. Finally, I revisit five RCTs in microcredit where average effects are not statistically significant and find evidence of both positive and negative treatment effects in household income. On average across studies, at least $13.6\%$ of households benefited, and $12.5\%$ were negatively affected.
@@ 1 Introduction
Important questions in impact evaluation cannot be answered by average treatment effects alone. What proportion of individuals are harmed by a treatment? Does a policy help many by a little? Or a few by a lot? The distribution of treatment effects carries crucial information with important equity and efficiency implications. For example, if a policy with positive average effect is found to harm a significant portion of the population, it may be crucial to prioritize mitigating the harm or reconsider its implementation. If a few individuals have large benefits, it may be important to understand why and how to efficiently target the policy recipients. However, learning the distribution of treatment effects is challenging because of the fundamental problem of causal inference: one can never observe individual counterfactuals. But one might be able to predict them.
@@
for any value of $\delta$, by leveraging covariate-adjustment and predictive models. The setting assumes a binary treatment $D$ and that the researcher has access to a representative sample of $Y(1)$ and $Y(0)$, the potential outcomes under treatment and no treatment, respectively, and pre-treatment covariates $X$. For example, this is often the setting in Randomized Controlled Trials (RCTs). Alternatively, one can rely on the selection on observables assumption with a known propensity score (Online Appendix OA.1). I denote $\theta$ by Distributional Treatment Effect (DTE) and the method by CAIDE (Covariate Adjustment for Inference on Distributional Effects). All results are pointwise in $\delta$.
Relative to previous literature, my contribution is to provide a new characterization of the identified set of $\theta$, from which I develop two novel methods for inference. The first is valid in finite samples and requires arguably weak conditions. For example, the conditions are satisfied if the data come from an RCT. To my knowledge, this is the first finite-sample inference result on the DTE that incorporates covariates. The second inference method is asymptotically valid and more powerful than the finite-sample approach. Both inference methods rely on estimating a covariate-adjustment term, and I show they are robust to its misspecification. If the covariate-adjustment term is correctly specified, the identified set is sharp. If it is misspecified, the identified set is still valid, although wider. As a consequence, the covariate-adjustment term can be estimated using any method, including machine learning algorithms.
My results can be contrasted with those in Fan and Park (2010), which characterized the sharp identified set of $\theta$ in the presence of covariates. Their approach to inference focuses on the case without covariates since it is challenging to implement inference based on their characterization of the identified set, as discussed in Section 2. In contrast, my characterization of the sharp identified set is more amenable to inference and satisfies the robustness property mentioned above.
@@
To understand my approach, consider a scenario where a researcher aims to predict $Y(0)$ using a function of covariates $s(X)$. An example of such a function is the conditional expectation, $s(x)=\mathbb{E}[Y(0)|X=x]$. Ideally, $s(X)$ would be close to $Y(0)$, and the distribution of $Y(1)-s(X)$ would approximate the distribution of treatment effects. However, $s(X)$ is generally an imperfect predictor for two main reasons. First, $Y(0)$ is typically random, meaning that even with many covariates, one can only approximate $Y(0)$ but not predict it perfectly. Second, nonparametric methods, including machine learning algorithms, are often biased and exhibit slow convergence rates, especially with many covariates. CAIDE is designed to address both of these challenges. In this example, the idea is that even if $s(X)$ is an imperfect predictor, examining the distribution of $Y(0)-s(X)$ (the prediction error) can help correct for the difference between $Y(1)-s(X)$ and the actual treatment effect. Intuitively, if $|Y(0)-s(X)|$ is small (e.g., less than 0.1), the error between $Y(1)-s(X)$ and the treatment effect is also small (at most $0.1$). In general, correcting for prediction error using the entire distribution of $Y(0)-s(X)$ provides more information than simply using an upper bound.
@@
The intuition above does not rely on the form of $s$. In fact, by combining the information from the distributions of both $Y(1)-s(X)$ and $Y(0)-s(X)$, any function $s$ can be used to learn $\theta$. However, when correcting for prediction error, different choices of $s$ can lead to different approximations of the distribution of treatment effects. CAIDE proceeds by estimating an optimal form of $s$, denoted by $s^{*}$ (defined in Theorem 1), which depends on the distributions of both potential outcomes conditional on covariates — that is, it is not the prediction of just one potential outcome. In this case, $s^{*}(X)$ can be interpreted as a covariate-adjustment term since the same function is subtracted from both $Y(1)$ and $Y(0)$.
@@
The fact that any function $s$ is valid motivates a sample-splitting strategy: one portion of the sample is used to estimate $s^{*}$ with $\hat{s}$ using any machine learning algorithm; in the remaining sample, I calculate a confidence interval for $\theta$ taking $\hat{s}$ as fixed. Once this function is fixed, estimation becomes a low-dimensional problem, similar to Fan and Park (2010), and valid inference can be achieved regardless of the statistical properties of $\hat{s}$. In fact, finite-sample inference on $\theta$ (Theorem 2) is valid when data come from an RCT, without requiring further regularity conditions. Moreover, cross-fitting, a more efficient way of sample-splitting, can be used for asymptotic inference under suitable conditions (Theorem 4). In both cases, if $\hat{s}$ is a poor estimate of $s^{*}$, the resulting confidence interval is wider but still valid compared to an estimator $\hat{s}$ that is close to $s^{*}$. I compare both approaches to inference in Section 5 through a Monte Carlo study.
@@
The rest of this paper is organized as follows. Section 1.1 reviews the related literature. Section 2 describes the setup and shows the main identification results of the paper. Section 3 provides finite-sample inference, and Section 4 shows asymptotic properties using cross-fitting. Section 5 presents a Monte Carlo study, and Section 6 the results of the application to microcredit. Finally, Section 7 concludes. Proofs not displayed in the main text are deferred to the Appendix.
@@ 1.1 Related Literature
Fan and Park (2010) also characterize sharp bounds on $\theta$ in the presence of pre-treatment covariates $X$, but they do not approach inference based on these bounds. It is challenging to provide inference from their characterization, as discussed in Section 2. In contrast, by incorporating covariates via a covariate-adjustment term, the characterization of the sharp bounds that I provide is amenable to inference.
@@
This paper also contributes to the literature on using machine learning for estimating heterogeneous treatment effects (Athey and Imbens, 2016; Chernozhukov et al., 2018; Wager and Athey, 2018; Athey et al., 2019; Yadlowsky et al., 2021; Athey et al., 2023; Chernozhukov et al., 2023). Most papers in this literature focus on estimating the point-identified conditional average treatment effect (CATE) and use the estimated CATE to predict individual treatment effects. CAIDE is complementary to these approaches. I focus on the partially-identified DTE $\theta$, which answers a different type of question. For example, if the CATE is zero for all values of $X$, many individuals could still be harmed, and many could benefit. In this case, the DTE could reveal, for example, the proportion of individuals who are harmed and who benefit.
This paper is also related to the literature on conformal prediction (Vovk et al., 2005; Lei et al., 2018). Lei and Candès (2021) propose a prediction interval to the individual treatment effect, i.e., an interval to cover the random variable $Y(1)-Y(0)$ for a specific individual with a certain probability. In contrast, I focus on the population parameter $\theta=P(Y(1)-Y(0)\leq\delta)$. Also related is the literature on covariate adjustment in the context of estimating the Average Treatment Effect (ATE) in RCTs (e.g., Lin, 2013, Wager et al., 2016, Rothe, 2018, Wu and Gagnon-Bartsch, 2018; review in Van Lancker et al., 2023; comparisons in Kahan et al., 2016, Tackney et al., 2023). Most of this literature focuses on adjusting outcomes for covariates to improve precision. I use covariate adjustment to improve identification (i.e., to get narrower bounds), not the power, of a different estimator, the DTE.
@@ 2 Setup and Identification via Covariate Adjustment
The second result is Theorem 1, which characterizes optimal forms of $s$ as to make the identified set in (2.3) sharp, that is, the narrowest possible interval that has the guarantee to contain $\theta$. As it turns out, in general, no single choice of $s$ makes both lower and upper bounds sharp; instead, each bound has a different function $s$. The sharp identified set in Theorem 1 is an alternative characterization to the one provided in Fan and Park (2010). Fan and Park (2010) do not address inference on $\theta$ in the presence of covariates. There are two main challenges for inference based on their identifying equation. First, if the conditional distributions of potential outcomes are misspecified, their characterization leads to invalid bounds, which may not contain $\theta$. The issue of misspecification is particularly relevant if the conditional distributions are estimated with bias, which is often the case when using nonparametric methods. The second is that the asymptotic distribution of plug-in estimators is not trivial to characterize, and it depends on the rate of convergence of the conditional distribution estimators. In contrast, misspecification or a slow rate of convergence of the covariate-adjustment term only affects the width of the bounds in (2.3), and the asymptotic distribution is easier to characterize (Theorem 3). Finally, in subsection 2.1 I discuss cases when $\theta$ is point-identified, which happens when covariates are informative enough to predict the counterfactuals accurately.
@@
Assumption A1(i) of bounded $\mathcal{Y}$ is technical, and is discussed in Remark 2. A1(ii) is the usual unconfoundedness assumption, and it is used for identification of the marginals of $Y(1)$ and $Y(0)$. It often holds, for example, in Randomized Controlled Trials (RCTs), where $D$ is randomized. Relaxing the assumption of constant propensity score is possible, and I propose two alternatives in Online Appendix OA.1. The first allows the propensity score to vary across groups of $X$, and the second allows it to vary arbitrarily with $X$ as long as $p(x)$ is known.
The starting point for CAIDE is the fact that for any function $s:\mathcal{X}\rightarrow\mathcal{Y}$, it holds that
@@
where $F_{j}(t)=P(Y(j)\leq t)$, $j\in\{0,1\}$. Hence, replacing the random variables $Y(1)$ and $Y(0)$ by $Y(1)-s(X)$ and $Y(0)-s(X)$, the bounds induced by $s$ contain $\theta$:
@@
Because of the equivalence in (2.1), and the validity of the Makarov bounds (2.2), (2.3) contains $\theta$ for any $s$. If covariates are not used, e.g. if $s(x)=0$, then (2.3) is equivalent to (2.2). Different functions $s$ lead to different bounds, with varying lengths. For specific functions, the induced bounds are sharp, that is, the smallest possible valid bounds given the information from the data. Fan and Park (2010) derived one characterization for these bounds:
@@
Functions $s$ that are different from the optimal $s^{*}_{L}(x)$ and $s^{*}_{U}(x)$ can lead to wider bounds, including wider than when $s(x)=0$. For example, consider a scenario where outcomes are constant, such as $Y(1)=1$ and $Y(0)=0$. Then, $\theta=P(Y(1)-Y(0)\leq 0)$ is point-identified and equal to zero. Let $s(X)=0$ for half of the population and $s(X)=1$ for the other half. Then, $Y(1)-s(X)\in\{0,1\}$ and $Y(0)-s(X)\in\{-1,0\}$, and the bounds induced by $s$ are $[0,1]$.

### v3

@@ Abstract
Important questions for impact evaluation require knowledge not only of average effects, but of the distribution of treatment effects. The inability to observe individual counterfactuals makes answering these empirical questions challenging. I propose an inference approach for points of the distribution of treatment effects by incorporating predicted counterfactuals through covariate adjustment. I provide finite-sample valid inference using sample-splitting, and asymptotically valid inference using cross-fitting, under arguably weak conditions. Revisiting five randomized controlled trials on microcredit that reported null average effects, I find important distributional impacts, with some individuals helped and others harmed by the increased credit access.
@@ 1 Introduction
Important questions in impact evaluation cannot be answered by average treatment effects alone. What proportion of individuals are harmed by a treatment? Does a policy help many by a little? Or a few by a lot? The distribution of treatment effects carries crucial information with important equity and efficiency implications. For example, if a policy with positive average effect is found to harm a significant portion of the population, it may be crucial to prioritize mitigating the harm or reconsider its implementation. If a few individuals have large benefits, it may be important to understand why and how to efficiently target the policy recipients. However, learning the distribution of treatment effects is challenging because of the fundamental problem of causal inference: one can never observe individual counterfactuals. But one may be able to predict them.
@@
for any value of $\delta$, by leveraging covariate-adjustment and predictive models. The setting assumes a binary treatment $D$ and that the researcher has access to a representative sample of $Y(1)$ and $Y(0)$, the potential outcomes under treatment and no treatment, respectively, and pre-treatment covariates $X$ [^1] In Online Appendix OA.1, I provide an alternative setting with a known propensity score. This is typically the setting in Randomized Controlled Trials (RCTs). I denote $\theta$ by Distributional Treatment Effect (DTE) and the method by CAIDE (Covariate Adjustment for Inference on Distributional Effects). All results are pointwise in $\delta$.
Relative to previous literature, my first contribution is to provide a new characterization of the identified set of $\theta$ which is more amenable to inference. Based on this result, I develop two novel methods for inference. The first is valid in finite samples and requires arguably weak conditions. For example, the conditions are satisfied if the data come from an RCT. To my knowledge, this is the first finite-sample inference result on the DTE that incorporates covariates. The second inference method is asymptotically valid and more powerful than the finite-sample approach. Both inference methods rely on estimating a covariate-adjustment term, and I show they are robust to its misspecification, in the sense that a misspecified term leads to a valid identified set, although wider. If the covariate-adjustment term is correctly specified, the identified set is sharp. As a consequence of this robustness, the covariate-adjustment term can be estimated using any method, including machine learning algorithms. The relevance of my contribution is demonstrated in an application to microcredit, where I find important distributional impacts, specifically statistically significant bounds on the proportion of individuals helped and hurt by increased credit access. While academics and policymakers have hypothesized and debated the theory underlying such impacts, their empirical estimation has been hampered by methodological limitations.
My results can be contrasted with those in Fan and Park (2010), who characterized the sharp identified set of $\theta$ in the presence of covariates. Their approach to inference focuses on the case without covariates since it is challenging to implement inference based on their characterization of the identified set, as discussed in Remark 1. In contrast, my characterization of the sharp identified set is more amenable to inference and satisfies the robustness property mentioned above.
@@
To understand my approach, consider a scenario where a researcher aims to predict $Y(0)$ using a function of covariates $s(X)$. An example of such a function is the conditional expectation, $s(x)=\mathbb{E}[Y(0)|X=x]$. Ideally, $s(X)$ would be close to $Y(0)$, and the distribution of $Y(1)-s(X)$ would approximate the distribution of treatment effects. However, $s(X)$ is generally an imperfect predictor for two main reasons. First, because $Y(0)$ is rarely a deterministic function of observable covariates, so one can typically only approximately predict it. Second, because $s(X)$ needs to be calculated from the data, and is subject to random sampling error. CAIDE is designed to address both of these challenges. In this example, the idea is that even if $s(X)$ is an imperfect predictor, examining the distribution of $Y(0)-s(X)$ (the prediction error) can help correct for the difference between $Y(1)-s(X)$ and the actual treatment effect. Intuitively, if $|Y(0)-s(X)|$ is small (e.g., less than 0.1), the error between $Y(1)-s(X)$ and the treatment effect is also small (at most $0.1$). In general, correcting for prediction error using the entire distribution of $Y(0)-s(X)$ provides more information than simply using an upper bound.
@@
The intuition above does not rely on the form of $s$. In fact, by combining the information from the distributions of both $Y(1)-s(X)$ and $Y(0)-s(X)$, any function $s$ can be used to learn $\theta$. However, when correcting for prediction error, different choices of $s$ can lead to different approximations of the distribution of treatment effects. CAIDE proceeds by estimating an optimal form of $s$, defined in Theorem 1, which depends on the distributions of both potential outcomes conditional on covariates — that is, it is not the prediction of just one potential outcome. In this case, the function $s$ can be interpreted as a covariate-adjustment term, since the same function is subtracted from both $Y(1)$ and $Y(0)$. In Section 2, I show that the optimal function is different for learning the lower or upper bound on $\theta$, and the procedure can be repeated with different functions $s_{L}^{*}$ and $s_{U}^{*}$ when one is interested in two-sided inference.
@@
The fact that any function $s$ is valid motivates a sample-splitting strategy: one portion of the sample is used to estimate $s^{*}$ ($s^{*}_{L}$ or $s^{*}_{U}$) with some estimator $\hat{s}$ using any machine learning algorithm; in the remaining sample, I calculate a confidence interval for $\theta$ treating $\hat{s}$ as fixed. Once this function is fixed, estimation becomes a low-dimensional problem, similar to Fan and Park (2010), and valid inference can be achieved regardless of the statistical properties of $\hat{s}$. In fact, finite-sample inference on $\theta$ (Theorem 2) is valid when data come from an RCT, without requiring further regularity conditions. Moreover, cross-fitting, a more efficient way of sample-splitting, can be used for asymptotic inference under suitable conditions (Theorem 4). In both cases, if $\hat{s}$ is a poor estimate of $s^{*}$, the resulting confidence interval is wider but still valid compared to an estimator $\hat{s}$ that is close to $s^{*}$. I compare both approaches to inference in Section 5 through a Monte Carlo study.
@@
The paper proceeds as follows. Section 1.1 reviews the related literature. Section 2 describes the setup and shows the main identification results of the paper. Section 3 provides finite-sample inference, and Section 4 asymptotic inference and exactness using cross-fitting. Section 5 presents a Monte Carlo study, and Section 6 the application to microcredit. Finally, Section 7 concludes. Proofs not displayed in the main text are deferred to the Appendix.
@@ 1.1 Related Literature
Fan and Park (2010) also characterize sharp bounds on $\theta$ in the presence of pre-treatment covariates, but they do not approach inference based on these bounds. It is challenging to provide inference from their characterization, as discussed in Remark 1. In contrast, by incorporating covariates via a covariate-adjustment term, the characterization of the sharp bounds that I provide is amenable to inference.
@@
This paper also contributes to the literature on using machine learning for estimating heterogeneous treatment effects (Athey and Imbens, 2016; Chernozhukov et al., 2018; Wager and Athey, 2018; Athey et al., 2019; Yadlowsky et al., 2021; Knaus et al., 2022; Athey et al., 2023; Chernozhukov et al., 2023). Most papers in this literature focus on estimating the point-identified conditional average treatment effect (CATE) and use the estimated CATE to predict individual treatment effects. CAIDE is complementary to these approaches. I focus on the partially-identified DTE $\theta$, which answers a different type of question. For example, if the CATE is zero for all values of $X$, many individuals could still be harmed, and many could benefit. In this case, the DTE could reveal, for example, the proportion of individuals who are harmed and who benefit.
This paper is also related to the literature on conformal prediction (Vovk et al., 2005; Lei et al., 2018). Lei and Candès (2021) propose a prediction interval to the individual treatment effect, i.e., an interval to cover the random variable $Y(1)-Y(0)$ for a specific individual with a certain probability. In contrast, I focus on the population parameter $\theta=P(Y(1)-Y(0)\leq\delta)$. Also related is the literature on covariate adjustment in the context of estimating the Average Treatment Effect (ATE) in RCTs (e.g., Lin, 2013, Wager et al., 2016, Rothe, 2018, Wu and Gagnon-Bartsch, 2018; review in Van Lancker et al., 2023; comparisons in Kahan et al., 2016, Tackney et al., 2023). Most of this literature focuses on adjusting outcomes for covariates to improve precision. I use covariate adjustment to improve identification (i.e., to get narrower bounds) of a different estimand, the DTE.
@@ 2 Setup and Identification via Covariate Adjustment
The second result is Theorem 1, which characterizes optimal forms of $s$ that make the identified set in (2.3) sharp, that is, the narrowest possible interval that has the guarantee to contain $\theta$. As it turns out, in general, no single choice of $s$ makes both lower and upper bounds sharp; instead, each bound requires a different function $s$, denoted $s^{*}_{L}$ and $s^{*}_{U}$ respectively for the lower and upper bounds. The sharp identified set in Theorem 1 is an alternative characterization to the one provided in Fan and Park (2010). I argue in Remark 1 that my representation leads to two major advantages for inference. Most importantly, I argue that inference using my characterization is tractable under arguably mild conditions (Theorems 2 and 4) when the bounds are estimated nonparametrically. In contrast, I show that it is hard to characterize the asymptotic distribution of the plug-in estimator based on Fan and Park (2010)’s characterization, even when using sample-splitting, when the bounds are estimated nonparametrically. Note that Fan and Park (2010) do not address inference on $\theta$ in the presence of covariates.
Finally, in subsection 2.1 I discuss cases when $\theta$ is point-identified, which happens when covariates are informative enough to predict the sign of the difference of counterfactuals.
@@
Assumption A1(i) of bounded $\mathcal{Y}$ is technical, and is discussed in Remark 2. A1(ii) is the usual unconfoundedness assumption, and it is used for identification of the marginals of $Y(1)$ and $Y(0)$. It typically holds, for example, in Randomized Controlled Trials (RCTs), where $D$ is randomized. The constant propensity score assumption facilitates exposition. In the asymptotic analysis, this assumption can be relaxed to allow for an arbitrary, known propensity score, as demonstrated in Online Appendix OA.1. Extending my results to accommodate unknown propensity scores that must be estimated from the data is an interesting avenue for future research.
The starting point for CAIDE is the fact that for any measurable function $s:\mathcal{X}\rightarrow\mathcal{Y}$, it holds that
@@
where $F_{j}(t)=P(Y(j)\leq t)$, $j\in\{0,1\}$. Hence, by replacing the random variables $Y(1)$ and $Y(0)$ by $Y(1)-s(X)$ and $Y(0)-s(X)$, the bounds induced by $s$ contain $\theta$:
@@
Because of the equivalence in (2.1), and the validity of the Makarov bounds (2.2), (2.3) contains $\theta$ for any $s$. If covariates are not used, e.g. if $s(x)=0$, then (2.3) is equivalent to (2.2). Different functions $s$ lead to different bounds, with varying lengths. For specific functions defined in Theorem 1, the induced bounds are sharp, that is, the smallest possible valid bounds given the information from the data. Fan and Park (2010) derived one characterization for these bounds:
@@
I argue that my representation of the identified set in Theorem 1 has two major advantages for inference compared to Fan and Park (2010)’s characterization in (2.5).
The first advantage is that Theorem 1 enables a tractable approach to inference under arguably mild conditions when the conditional distributions are estimated nonparametrically (Theorems 2 and 4). In contrast, deriving the asymptotic distribution of the plug-in estimator based on Fan and Park (2010)’s characterization is challenging, even when using sample-splitting, when the conditional distributions are estimated nonparametrically. I show in Appendix A.3 that this asymptotic distribution depends on the specific estimation method used for the conditional distributions, and the convergence rate of the plug-in estimator depends on how quickly they are learned. It is difficult to characterize this asymptotic distribution without an assumption on the convergence rates of the conditional distribution estimators. However, such assumption often does not hold when the dataset includes more than a few covariates and the conditional distributions are estimated nonparametrically.
The second advantage of my identifying equation is that the bounds in (2.3) are robust to misspecification of the covariate-adjustment terms when they are learned parametrically. That is, if $(s^{*}_{L},s^{*}_{U})$ are estimated inconsistently, the limit bound may be wider than the sharp bounds, but they always contain $\theta$. As a consequence, the inference results established in Theorems 2 and 4 are valid regardless of whether $(s^{*}_{L},s^{*}_{U})$ are estimated nonparametrically or with a potentially misspecified parametric model. In contrast, the plug-in estimator based on Fan and Park (2010)’s characterization can converge to bounds narrower than the sharp, and thus not contain $\theta$, if the conditional distributions are estimated with a misspecified parametric model.
Moreover, this robustness property is also advantageous when $(s^{*}_{L},s^{*}_{U})$ are estimated consistently, since noisy estimates of the covariate-adjustment terms translate into estimating bounds that are wider. In contrast, the plug-in estimator based on Fan and Park (2010)’s characterization does not have this property. This robustness property is relevant both for the finite-sample valid inference approach of Theorem 2, which relies only on an RCT assumption, and the large-sample approach of Theorem 4.
@@
Remark 3.
In principle, functions $s$ that differ from the optimal $s^{*}_{L}(x)$ and $s^{*}_{U}(x)$ can lead to wider bounds, potentially even wider than those obtained using $s(x)=0$. This phenomenon, however, is not of concern when $s^{*}_{L}(x)$ and $s^{*}_{U}(x)$ are estimated nonparametrically, since they yield the sharp bounds, which are by construction (weakly) narrower than those using $s(x)=0$.